Suppose that two firms are producers of spring water, which can be obtained at zero cost. The marginal revenue for their combined output is MR =10-2Q where MR is marginal revenue and Q is the number of litres of spring water sold by both firms per hour. If the 2 producers act according to the oligopoly model, how much will their combined output be? How much is the market price of water per litre?
1
Expert's answer
2020-10-25T19:16:10-0400

profit maximize condition will be achieved

at MR = MC

"10-2a = 0"

"2a = 10"

"a = 5"

total revenue = price "\times" quantity

TR = PQ

TR = "\int" MR "d\theta"

TR "= \int""(10-2a)d\theta"

TR = "10a-a^2"

"p= \frac{TR}{a}"

"= \frac{(10-a^2)}{a}"

"= 10-a"

"p= 10-5"

"=5"

market price of water per liter = $5